### v1

# Mitigating the Stability-Plasticity Dilemma in Adaptive Train Scheduling with Curriculum-Driven Continual DQN Expansion

Abstract
A continual learning agent builds on previous experiences to develop increasingly complex behaviors by adapting to non-stationary and dynamic environments while preserving previously acquired knowledge. However, scaling these systems presents significant challenges, particularly in balancing the preservation of previous policies with the adaptation of new ones to current environments. This balance, known as the stability-plasticity dilemma, is especially pronounced in complex multi-agent domains such as the train scheduling problem, where environmental and agent behaviors are constantly changing, and the search space is vast. In this work, we propose addressing these challenges in the train scheduling problem using curriculum learning. We design a curriculum with adjacent skills that build on each other to improve generalization performance. Introducing a curriculum with distinct tasks introduces non-stationarity, which we address by proposing a new algorithm: Continual Deep Q-Network (DQN) Expansion (CDE). Our approach dynamically generates and adjusts Q-function subspaces to handle environmental changes and task requirements. CDE mitigates catastrophic forgetting through EWC while ensuring high plasticity using adaptive rational activation functions. Experimental results demonstrate significant improvements in learning efficiency and adaptability compared to RL baselines and other adapted methods for continual learning, highlighting the potential of our method in managing the stability-plasticity dilemma in the adaptive train scheduling setting.
Introduction
Continual learning agents, capable of acquiring new knowledge and skills over time, are essential for environments that constantly change (Abel et al. 2024; Khetarpal et al. 2022). One central challenge for continual learning systems is balancing the retention of previously learned knowledge with the acquisition of new knowledge—a concept known as the stability-plasticity dilemma (Schwarz et al. 2018). This dilemma is relevant in many real life application domains such as the train scheduling problem (TSP), where environmental conditions (i.e routes, schedules) and agent behaviors (i.e train characteristics) are constantly changing, and the search space is vast. TSP is inherently complex. Even in simplified scenarios, agents often require considerable time to converge on optimal solutions (Toth and Vigo 2002).
One approach to address the complexity of TSP is curriculum learning, which introduces a sequence of environments that gradually increase in difficulty, allowing the agent to build upon previously acquired skills (Narvekar et al. 2020). This approach mirrors how humans learn, starting with simpler concepts and progressively tackling more complex ones. However, curriculum learning introduces its own set of challenges due to the non-stationary training environment.
Various continual reinforcement learning (CRL) approaches have been proposed to mitigate catastrophic forgetting, improve plasticity, facilitate transfer learning across new tasks, and scale effectively to handle multiple tasks. Current methods can be categorized into fixed-size approaches, which do not change the size of the network for new tasks (i.e., regularization-based approaches or optimization-based methods), and network expansion approaches. While network expansion approaches are more likely to achieve better performance, they may suffer from scaling issues as the number of tasks increases (Wang et al. 2023).
To address this, recent studies in the mode connectivity literature have investigated whether multitask and continual solutions are connected by a manifold of low error. Results show that different optima obtained by gradient-based optimization methods are connected by simple paths of non-increasing loss (Mirzadeh et al. 2020). Inspired by this, various works have demonstrated how to efficiently build simplicial complexes for fast ensembling and adaptively grow the subspace as the system encounters new tasks. This can be leveraged to develop neural systems that scale better while maintaining good performance across multiple tasks (Benton et al. 2021; Gaya et al. 2022).
While previous works focused on continual learning with distinct tasks, a curriculum does not involve distinct tasks with clear boundaries but rather a series of adjacent skills that build on top of each other, such as pathfinding, dealing with unexpected malfunctions, or avoiding deadlocks. This introduces a unique type of non-stationarity, where agents are learning interrelated skills. The challenge lies in effectively sequencing these skills to maximize learning efficiency and performance while also addressing the stability-plasticity dilemma for the learning agents.
In this work, we propose a well-structured curriculum for the train scheduling problem, decomposing it into various skills like pathfinding, malfunction handling, and deadlock avoidance. Furthermore, we investigate different approaches to leverage this curriculum and introduce the Continual Deep Q-Network Expansion (CDE) algorithm. The core idea of our CDE approach is to dynamically expand, adjust, and prune Q-function subspaces based on current task changes in the environment. The decision-making process within CDE involves determining whether a pre-existing Q-function subspace is effective for a new task and thus adapting it, or if a new subspace needs to be created for faster learning. The pre-existing Q-function subspaces are adapted to new tasks with elastic weight consolidation to avoid catastrophic forgetting. while the newly created subspaces are trained with adaptive rational activation functions for higher plasticity. In the pruning step, we retain only the best Q-networks to maintain sublinear scaling. Experimental results demonstrate significant improvements compared to other approaches, highlighting the potential of CDE to balance stability and plasticity, leading to more efficient and adaptable learning [^1] A link to the code will be added soon..
Figure 1: Diagram illustrating the Continual DQN Expansion Algorithm. When a new task or a change in the environment occurs, a new subspace with anchor $\theta$ is initialized. The new Q-function is trained with adaptive rational activation functions ($\theta^{HP}_{i}$) to enhance plasticity and achieve faster convergence on the new task. Simultaneously, the previous subspaces are further adapted on the new task by training Q-functions ($\theta^{CF}_{i}$) using elastic weight consolidation (EWC) regularization to mitigate catastrophic forgetting on previous tasks. Depending on the performance of different subspaces, the subspace set is either extended or pruned.
Related Work
Continual Learning Traditional RL methods, which assume static environments, often lead to models prone to overfitting and lacking adaptability to new, varying contexts. Continual RL (CRL) aims to develop adaptive models that learn from a continuous data stream and evolve across multiple tasks by leveraging accumulated knowledge, thereby overcoming the limitations of static-task assumptions (Abel et al. 2024).
Addressing Catastrophic Forgetting: Efforts such as those by Rusu et al. (2016), Cheung et al. (2019), and Wortsman et al. (2020) aim to preserve model parameters from previously learned tasks but often do not scale well with an increasing number of tasks mainly due to substantial computational and memory requirements. Other strategies, like maintaining a memory buffer of prior experiences (Lopez-Paz and Ranzato 2017; Isele and Cosgun 2018; Caccia et al. 2020), also face scalability issues as memory costs increase with task complexity and number. Methods enhancing transfer to new tasks, such as fine-tuning, are scalable and adaptable but prone to catastrophic forgetting. Elastic weight consolidation (Kirkpatrick et al. 2017) addresses this by limiting weight modifications, though it may reduce network’s plasticity as the number of tasks increase. Knowledge distillation techniques, as discussed in (Rusu et al. 2015; Li and Hoiem 2017; Rolnick et al. 2019), and methods leveraging commonalities across tasks (Xu and Zhu 2018; Pasunuru and Bansal 2019), have been utilized to improve CRL performance. Meta-learning and generative models also enhance adaptability and learning efficiency (Javed and White 2019; Beaulieu et al. 2020; Shin et al. 2017). However, practical application challenges persist, as indicated by (Lee et al. 2020), who propose an iteratively expanding network model primarily for supervised learning without distinct task boundaries.
Addressing Neural Plasticity: While addressing catastrophic forgetting is essential for non-stationary agents, extensive neural plasticity is required due to inherent distribution shifts during agent training, such as input and output drifts (Mnih et al. 2015). These shifts are exemplified by various environment dynamics as agents adapt to changing states and rewards. Recent studies have explored approaches to increase plasticity and allow faster adaptation to changing environments, such as concatenated ReLU (Abbas et al. 2023) and the introduction of new trainable weights (Nikishin et al. 2024). Additionally, techniques like continual backpropagation (Dohare et al. 2023) and dynamic hyperparameter and activation function adaptation (Loni et al. 2023; Delfosse et al. 2021) highlight the potential of dynamically alter the weight optimization landscape,which can aid continual adaptation and thus paving the way for more robust RL systems.
Curriculum Learning Curriculum learning is a learning paradigm where training data is organized in a progressively challenging manner, introducing concepts gradually to improve learning outcomes (Bengio et al. 2009). Curriculum learning is applied across various fields demonstrating improved sample efficiency and better convergence of machine learning algorithms (Soviany et al. 2022). Curriculum learning not only helps in achieving better performance on evaluation data but also in handling noisy datasets by starting with cleaner data and gradually introducing noisier parts (Wang, Chen, and Zhu 2021).
In reinforcement learning, curriculum learning has the potential to enhance learning efficiency and performance(Narvekar et al. 2020). Early applications of curriculum learning in RL were primarily manual, such as the work by Stanley and Miikkulainen (2005), where domain experts designed curricula for video game agents. Recently, there has been a shift towards automating curriculum generation. For instance, Narvekar and Stone (2019) proposed learning a curriculum policy using a Curriculum Markov Decision Process (CMDP), which maps the learning progress to the subsequent tasks, showing that intermediate tasks do not always need to be learned to convergence. Other methods, such as those by Foglino, Coletto Christakou, and Leonetti (2019) and Foglino et al. (2019), frame curriculum sequencing as a combinatorial optimization problem, employing metaheuristic algorithms like beam search and genetic algorithms to identify optimal task sequences.
Despite these advancements, several challenges persist in curriculum learning for RL. One significant issue is the computational expense associated with learning a curriculum policy independently for each agent and task, as highlighted by Narvekar and Stone (2019). Another challenge is catastrophic forgetting, where knowledge from previous tasks is lost when learning new tasks. Additionally, when state and action spaces differ between tasks, transferring high-level knowledge such as partial policies or options becomes crucial, as explored by Yang and Asada (1996) and Zimmer et al. (2018).
Train Scheduling Problem
Mathematical Formulation
TSP is an NP-Hard optimization and logistics challenge focused on the optimal assignment of trains to routes and their scheduling across a railway network while adhering to diverse constraints (Caprara, Fischetti, and Toth 2002). It requires dynamic replanning to adapt to unforeseen disruptions effectively. This issue integrates aspects of Multi-Agent Pathfinding and the Vehicle Rescheduling Problem, combining approaches from both fields to enhance the efficacy of train scheduling strategies (Cai and Goh 1994).
Consider a network of yards indexed from 00 to $n$. Each train $T_{k}$ is assigned a route $(i,j)$ where $i$ and $j$ are yards in the network. Trains have varying speeds and malfunction rates. In the case of a malfunction, a train is halted for a certain number of time steps.
Given $K$ (number of trains), $N=\{1,\ldots,n\}$ (set of yards), $S_{k}$ (stop-schedule for train $k$), $v_{k}$ (speed of train $k$), $m_{k}$ (malfunction rate of train $k$), $h_{k}$ (time steps train $k$ is halted in case of a malfunction), $d_{ij}$ (distance between yards $i$ and $j$), $x_{ij}^{k}$ (binary variable indicating if train $k$ travels directly from yard $i$ to yard $j$), the objective is to minimize the total travel time for all trains, including delays due to malfunctions. The mathematical formulation of the TSP can be represented as follows:
$\min\sum_{k=1}^{K}\sum_{(i,j)\in S_{k}}\left(\frac{d_{ij}}{v_{k}}+m_{k}h_{k}\right)x_{ij}^{k}$ (1)
subject to:
$\sum_{j\in N}x_{ij}^{k}=1\quad\forall k=1,\ldots,K,\,\forall i\in N$
$\sum_{i\in N}x_{ij}^{k}=1\quad\forall k=1,\ldots,K,\,\forall j\in N$
$\sum_{i\in N}x_{ij}^{k}\leq 1\quad\forall k=1,\ldots,K,\,\forall j\in N$
The first two constraints ensure that each train must have a complete route from a source yard to a destination yard. The third constraint ensures that a train cannot visit a yard more than once.
Traditionally, Operations Research (OR) techniques are employed to tackle the TSP, leveraging mathematical modeling and optimization to systematically improve train path scheduling (Lusby et al. 2011; Dong et al. 2020). These methods excel in breaking down the TSP into a graph optimization problem, aiming to identify non-intersecting routes for trains in time and space.
However, the reliance on OR comes with limitations, particularly in its flexibility and responsiveness to unexpected disruptions, such as weather conditions or infrastructure or train failures. Additionally, as railway networks expand and the number of operational trains increases, the computational demands of OR methods surge, leading to significant challenges in maintaining optimal schedules, especially under conditions requiring frequent rescheduling (Khadilkar 2018).
Flatland Simulator
The Flatland Simulator is two-dimensional simplified grid environment that allows for faster and diverse experimentation in TSP. Flatland does not only reduce the complexity of the full physical simulation, but also provides an easy-to-use interface to test novel approaches for the TSP (Mohanty et al. 2020).
Agents are tasked with efficiently navigating trains from start to finish, with the complexity of tasks escalating as the grid expands and more trains are added.
The score is computed based on the successful routing of trains to their destinations and the completion of scenarios, where all trains within an environment reach their respective endpoints, testing their strategies for effective train scheduling and rescheduling under varying conditions.
Curriculum Design for the Train Scheduling Problem
To construct an effective curriculum, it is essential to consider task generation, sequencing, and transfer learning. A well-designed curriculum improves both accuracy and speed by breaking down the train scheduling problem into distinct subproblems enabling better performance and faster convergence.
In designing our curriculum for the TSP, we focused on addressing four key challenges integral to efficient rail system management:
  • Pathfinding: An agent is expected to navigate the network efficiently from start to destination without unnecessary detours, recognizing that optimal pathfinding is foundational to handling complex scheduling scenarios.
  • Train-Speed Differences: An agent should also address conflicts arising from varying train speeds. Slower trains can hinder the progress of faster ones, necessitating strategic planning to mitigate such issues.
  • Malfunctions: An agent should be able to handle random malfunctions, necessitating adaptive responses to avoid delays.
  • Deadlocks: It is critical that an agent efficiently handles various types of deadlocks (common, cyclic, and track-caused; see Appendix for visualization and further details). An agent should efficiently plan and avoid paths that might lead to these scenarios, even at the cost of increased travel time for some trains.
These challenges often occur in combination, requiring agents to make decisions that minimize the impact of such scenarios on overall system efficiency.
Based on these challenges, we design a curriculum tailored to address specific train scheduling problems (TSP) by alternating between three specialized environment types focused on pathfinding, malfunctions and train-speed differences, and deadlocks (see Figure LABEL:fig:main).
  • Pathfinding starts with a simple 4x4 environment focusing on single-agent navigation, gradually increasing in size to enhance the agent’s goal-reaching capabilities.
  • Malfunctions and train-speed differences are taught in a unified setting due to their overlapping problem-solving strategies, emphasizing decision-making between following closely and opting for longer, less congested routes.
  • Deadlock avoidance focuses on the most common deadlock scenarios, gradually increasing complexity by adding more agents and reducing number of switches.
In this work, we sequence the tasks in the following way: Pathfinding, malfunctions and train-speed adjustments, and deadlock avoidance. We posit that starting with pathfinding is the most sensible choice, as all other tasks require an efficient pathfinding strategy. We leave deadlock avoidance for last, as it requires the agent to master the skills of pathfinding and handling malfunctions and various train speeds.
Furthermore, we consider two versions of our curriculum: with and without rehearsal. Rehearsal involves revisiting earlier tasks to reinforce learning, aiming to examine its impact on knowledge retention and application across varying complexities. The curriculum’s structure is designed to equally distribute learning across different environments. See the Appendix for detailed parameters of the curriculum.
Algorithm 1 Continual DQN Expansion (CDE)
0:  $\theta_{1},\ldots,\theta_{j}$ (previous Q- Approximators), $\epsilon$ (threshold), New Task $k$,
1:  Initialize replay buffer $R_{k}\leftarrow\emptyset$
2:  Initialize: $\theta_{j+1}\leftarrow\frac{1}{j}\sum_{i=1}^{j}\theta_{i}$ with learnable Rational Activation Functions
3:  for  Episode $e=1$$E$ do
4:     Initialize a random process $\mathcal{N}$ for action exploration
5:     Receive initial observation state $s_{1}\sim p_{K}(s_{1})$ for the new task $k$.
6:     for $t=1$$T_{k}$ do
7:        Select a random index $i\in\{0,\ldots,j+1\}$
8:        Select action $a_{t}=\mu(s_{t}|\theta^{\mu}_{i})+\mathcal{N}_{t}$
9:        Execute $a_{t}$ and observe $r_{t}$ and $s_{t+1}$
10:        Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $R_{k}$
11:     end for
12:     if Weight Update Step then
13:        Update The Q-Network with weights $\theta_{j+1}$ and their adaptive activation functions.
14:        Update The previous Q-Networks $\theta_{1},\ldots,\theta_{j}$ with Loss $L(\theta_{i})=L_{B}(\theta_{i})+\sum_{k}\frac{\lambda}{2}F_{i}(\theta_{i}-\theta_{i}^{*})^{2}$ where $\theta_{i}^{*}$ are the weights of the models trained on the previous task.
15:     end if
16:  end for
17:  if $W_{\phi}(\theta_{\text{j+1}})>\epsilon\cdot W_{\phi}(\theta_{i})$ for all $i\in\{1,\ldots,j\}$ then
18:     find $k=\arg\min_{i\in\{1,\ldots,j\}}W_{\phi}(\theta_{i})$
19:     return $\theta_{1},\ldots,\theta_{k-1},\theta_{k+1},\ldots,\theta_{j},\theta_{j+1}$;
20:  else
21:     return $\theta_{1},\ldots,\theta_{j}$;
22:  end if
Continual DQN Expansion Algorithm
We propose a novel approach for adaptively constructing Q-function subspaces to address the challenges of non-stationary environments. Our Continual Deep Q-Network (DQN) Expansion (CDE) approach dynamically generates and adjusts Q- function subspaces, denoted as $\Theta_{1},\ldots,\Theta_{N}$, based on environmental changes and task requirements. This method not only aims to mitigate catastrophic forgetting through selective expansion and the use of an elastic weight consolidation regularized loss but also seeks to enhance the agent’s plasticity by leveraging rational adaptive activation function for faster adaptability to new tasks.
In the subsequent sections, we provide a brief overview of elastic weight consolidation (EWC) and padé activation units (PAU). Following this, we explain our methodology, detailing the integration of our proposed subspace expansion technique with EWC and PAU.
Formalisation of Continual Reinforcement Learning Problem
A continual reinforcement learning (RL) scenario is characterized by a sequence of $N$ tasks or instances. Each task $k$ is described through a Markov Decision Process (MDP) $M_{k}=\langle S_{k},A_{k},T_{k},r_{k},\gamma\rangle$, which encompasses a set of states $S_{k}$, a set of actions $A_{k}$, a transition function $T_{k}:S_{k}\times A_{k}\rightarrow\mathcal{P}(S_{k})$, and a reward function $r_{k}:S_{k}\times A_{k}\rightarrow\mathbb{R}$. We assume in our case that all tasks share identical state and action spaces.
A global Q-function $Q:[1..N]\times\mathcal{Z}\rightarrow(S\times A\rightarrow\mathbb{R})$ is defined. This function accepts a task identifier $k$ and a sequence of tasks $\mathcal{Z}$ as inputs, yielding a Q-value representing the expected future rewards for interacting with the environment of $k^{\text{th}}$ task. The output is a real-valued estimate of the expected return from taking action $A$ in state $S$.
For each new task, we reset the replay buffer. Meaning that upon transitioning to task $k+1$, the transitions from the previous MDP $M_{i}$ are no longer accessible to the agent.
Elastic Weight Consolidation
The main idea behind behind EWC is that given a sequence of tasks, after learning task $k$, the parameters $\theta$ of the neural network are adjusted not only based on the current task’s loss but also considering the importance of these parameters to the previously learned tasks $11,\ldots,k-$ 1. EWC introduces a regularization term to the loss function (Kirkpatrick et al. 2017). The regularization term is computed based on the Fisher information matrix $F$ from the previous tasks, which estimates the importance of each parameter to the tasks already learned.
This term effectively constrains the parameter updates, favoring updates that have minimal impact on the performance of past tasks. Thus, EWC allows the neural network to learn new tasks while preserving performance on previously learned tasks by modulating the flexibility of its weights based on their importance to past learning.
The loss function $L(\theta)$ that we minimize in EWC is defined as:
$L(\theta)=L_{B}(\theta)+\sum_{j}\frac{\lambda}{2}F_{j}(\theta_{i}-\theta_{j,k}^{*})^{2}$ (2)
where $L_{k}(\theta)$ is the loss for task $k$ only, $\lambda$ sets the importance of the old task relative to the new one. $F_{j}(\theta)$ is a regularization term that penalizes changes to parameters that are critical for the performance on task $j$.
When moving to the next task, EWC aims to keep the network parameters close to the learned parameters of previous tasks. This can be enforced either with two separate penalties or as one by noting that the sum of two quadratic penalties is itself a quadratic penalty.
Padé Activation Units
Padé activation units (PAUs) enhance the adaptability and efficacy of neural networks in continual reinforcement learning environments (Delfosse et al. 2021). PAUs, defined as the ratio of two polynomial functions, offer a flexible and learnable activation mechanism that can dynamically adjust to varying task complexities. This adaptability is particularly beneficial in continual RL settings, where the agent encounters a sequence of tasks, each potentially differing in its dynamics and objectives.
In a continual RL setting, employing PAUs allows for the modulation of the activation functions based on the specific requirements of each task, thus improving the plasticity of neural networks. Formally, a padé activation function can be represented as:
$R(x)=\frac{\sum_{j=0}^{m}a_{j}x^{j}}{1+\sum_{k=1}^{n}b_{k}x^{k}}$ (3)
where $a_{j}$ and $b_{k}$ are the learnable parameters that determine the shape of the activation function. The adaptability of PAUs lies in their ability to represent a wide range of nonlinear relationships through the adjustment of these parameters, thereby enabling a neural network to maintain robust performance across a continuum of RL tasks and non-stationary environments.
Subspace Expansion of Q-Functions
We introduce the Continual Deep Q-Network (DQN) Expansion (CDE) approach, a novel method designed to adaptively construct Q-function subspaces for continual reinforcement learning. Our approach facilitates sequential task learning by an agent. The pseudo-code is provided in Algorithm 1.
The CDE algorithm generates a sequence of subspaces, denoted as $\Theta_{1},\ldots,\Theta_{N}$, each corresponding to a new change in the agent’s environment. The dynamic expansion process involves the integration of new anchors $\theta$ function, defined by a neural network. A new anchor is added only if it demonstrates a superior performance over the previous anchors within the existing subspace, ensuring sublinear expansion of the number of anchors. Our expansion is limited to exactly N networks as we remove the worst-performing anchors each time we add a new ones, if we already have N anchors. N is a hyperparamter that depends on the continual learning task. In our application, we set $N=2$. Higher N values likely leads to better performance but also cause higher computational cost for training.
CDE maintains the current anchors set composed of the weight vectors $\{\theta_{1},\ldots,\theta_{N}\}$ that delineate the optimal Q-functions for previously encountered tasks.
Upon the introduction of a new task, denoted as $k+1$, the models creates a preliminary new subspace $\Theta_{j+1}$. We initialize the weights of the new Q-function network as follows:
$\theta_{j+1}=\frac{1}{j}\sum_{i=1}^{j}\omega_{i}\cdot\theta_{i}$
where $\omega_{i}$ is the coefficient of each previous anchor. To mitigate catastrophic forgetting of previous tasks, the weights $\{\theta_{1},\ldots,\theta_{N}\}$ of the previous subspaces are refined using an EWC regularized loss. The anchor $\theta_{j+1}$ for the new subspace is optimized with rational adaptive activations to enhance plasticity and accelerate the adaptation process.
The decision-making process within CDE involves either identifying an effective pre-existing Q-function for the new task within the current subspace or creating a new subspace with a more adequate anchor $\theta_{j+1}$ for the new task. If the new subspace achieves higher performance and we have already reached the maximum subspace size, we remove the worst-performing subspace in the pruning step.
The determination to incorporate the new anchor is based on a assessment of the optimal Q-functions for task $k$ across both the newly formed and the previous subspaces.
Our approach can be summarized in three main steps: (1) expanding the current subspace by introducing a new anchor, (2) training and adjusting all anchors for the current task, and (3) evaluating the efficacy of the new anchor relative to the optimal anchor function within the existing subspace.
Based on this comparison, CDE determines whether the subspace should be further extended (i.e., adding a new anchor to the subspace set) or reverted to its previous configuration (keeping the subspace set as it is) or pruned (adding a new anchor and removing the worst-performing anchor). This ensures a balanced adaptation to new tasks while preserving previously acquired knowledge.
Experiments
Training Curricula
The standard approach to training reinforcement learning algorithms for the train scheduling problem involves directly training the agents in the target environment without any preliminary steps. For the Flatland scenario, this meant using a sparse line generator to create 64x64 environments with 14 agents. While this method is straightforward due to the uniformity of training and evaluation settings, the vast search space makes generalization difficult and leads to slower convergence.
We train and evaluate various reinforcement learning algorithms under the following learning conditions: (1) no curriculum, (2) dynamic or naive environment curriculum, and (3) our proposed custom curriculum.
The dynamic environment curriculum begins with less complex environments and progressively increases in difficulty by enlarging the environment size and agent count. This gradual introduction of complexity helps agents adapt to more challenging scenarios without the abrupt difficulty spike seen in the no curriculum setting. In contrast, our proposed custom curriculum addresses specific TSP challenges by alternating between three specialized environment types focused on pathfinding, malfunctions and train-speed differences, and deadlocks.
Experimental Setting
To evaluate the performance of different algorithms, we utilize a simulated environment that closely replicates a realistic and complex railway system. Each algorithm is tested without undergoing a prior adaptation phase on the test environment, allowing us to focus on their ability to generalize and apply learned skills in unfamiliar settings.
We use score and completion rate as metrics to evaluate the experiments. The score is the combined reward all agents receive at the end of an episode divided by the number of agents. The completion rate is the percentage of trains reaching their destination before the simulation ends. It is important to mention that a better score does not always correlate with a higher completion rate. This is because the score represents how close an agent got to its destination. In most cases, substantial improvements in the completion rate are required to achieve relatively small improvements in the score. Furthermore, the variance of the completion rate between runs is usually high, as observed in our experimental results. For these reasons, we report both the reward and completion rate in our experimental analysis.
Baselines
Our baselines consist of an Advantage Actor-Critic (A2C), Proximal Policy Optimization (PPO) , and Deep Q-Network (DQN) algorithm (Schulman et al. 2017; Mnih et al. 2015). Each algorithm utilizes two hidden layers with 1024 units each, a replay buffer of size 1 million a batch size of $128$, and a learning rate of $5\cdot 10^{-5}$. We use a simulated annealing exploration strategy.
For the rational padé activation units (PAU), we use values of 5 for $m$ and 4 for $n$, initializing with the shape of a Rectified Linear Unit. For EWC, we set the value of $\lambda$ to 0.5, determining how much EWC can hinder the further adaptation of certain neurons. Further implementation details are included in the appendix.
Results and Discussion
Method No Curric. Naive Curric. Custom Curric. Custom Curric. (Rehear.)
Score Completion Score Completion Score Completion Score Completion
A2C 0.34 ± 0.04 0.06 ± 0.10 $0.34\pm 0.04$ $0.07\pm 0.10$ $0.35\pm 0.04$ $0.05\pm 0.08$ $0.34\pm 0.04$ $0.09\pm 0.06$
PPO 0.34 ± 0.04 0.06 ± 0.09 $0.35\pm 0.04$ $0.06\pm 0.09$ $0.35\pm 0.04$ $0.06\pm 0.09$ $0.34\pm 0.04$ $0.06\pm 0.10$
DQN 0.35 ± 0.04 0.04 ± 0.09 $0.35\pm 0.04$ $0.08\pm 0.11$ $0.34\pm 0.04$ $0.10\pm 0.12$ $\mathbf{0.30\pm 0.04}$ $\mathbf{0.19\pm 0.06}$
Table 1: Performance comparison of different baselines trained with a range of curricula. The score (lower is better $\downarrow$) and completion rate (higher is better $\uparrow$) are reported. The best score and completion rate are highlighted in bold. We see that the custom curriculum with decomposed tasks (see Figure LABEL:fig:main) achieves the highest performance. Further improvements are observed for the custom curriculum with rehearsal in the case DQN.
In Table 1, we compare the performance of various standard RL algorithms trained using different curricula on the evaluation environment. The table illustrates that incorporating a curriculum enhances the performance of all baselines in terms of scores and completions. For instance, DQN achieves a score of $0.4\pm 0.03$ without a curriculum and a score of $0.35\pm 0.04$ with one of the curricula (lower scores are better). Specifically, our proposed curriculum improves DQN scores and completions compared to the naive curriculum, indicating that a more structured curriculum can be helpful. However, the improvements are less pronounced for both PPO and A2C, remaining within the margin of error. These results highlight the importance of pairing curricula with algorithms tailored to the application domain.
Furthermore, we observe that rehearsal significantly improves DQN’s performance, nearly doubling the completion rate from $0.10$ with the proposed custom curriculum to $0.19$ with the proposed custom curriculum plus rehearsal. This underscores the need to address catastrophic forgetting to learn a more generalizable strategy, which we aim to achieve with our proposed method, CDE. In the proposed curriculum, specific skills are learned in each new environment, so it is crucial to ensure the agent retains previous skills. For example, pathfinding skills are essential for handling varying train speeds, illustrating how tasks build on each other. Additional results in the appendix demonstrate how task ordering impacts generalization performance, emphasizing the need for domain understanding and expertise in creating meaningful curricula.
In Table 2, we extend DQN with different techniques and compare the generalization performance when training using our custom curricula. We observe that increasing plasticity (DQN +PAU) or mitigating catastrophic forgetting (DQN +EWC) helps achieve more generalizable minima compared to a simple DQN. However, both methods perform worse than simple rehearsal ($0.19$ completion for DQN +EWC and $0.12$ for DQN +PAU compared to $0.19$ for DQN with rehearsal). Additionally, in this scenario, addressing catastrophic forgetting is more beneficial than enhancing plasticity. This is because rational activations are most advantageous when the neural network begins to lose plasticity after a large number of tasks (e.g., more than 10 tasks). In contrast, our curriculum involves only three tasks, each comprising four different environment sizes.
A naive combination of both approaches for plasticity (PAU) and stability (EWC) results in performance degradation for DQN +EWC+PAU compared to just DQN +EWC. This can be explained by the opposing effects of rational activations and elastic weight consolidation, where changing activations render the Fisher matrix weights for EWC obsolete.
Instead, our method (CDE) ensures the network remains both plastic and stable depending on the current task. CDE achieves the best performance, with almost a 50% improvement in the completion rate compared to DQN with rehearsal, even without using curriculum rehearsal during training of CDE. Our approach balances plasticity and stability by expanding the network in a targeted manner based on the current task and the performance of pre-existing subspaces. Further results in the appendix show that CDE switches between more stable and more plastic networks depending on the task requirements, is less affected by the task ordering and achieves better generalization, and thus effectively mitigating the stability-plasticity dilemma in a curriculum learning setting.
Method Score Completion Rate
DQN 0.34 ± 0.04 0.10 ± 0.12
DQN w/ Rehear. 0.30 ± 0.04 0.19 ± 0.17
DQN +EWC 0.31 ± 0.04 0.19 ± 0.18
DQN +PAU 0.33 ± 0.04 0.12 ± 0.14
DQN +PAU+EWC 0.32 ± 0.04 0.18 ± 0.16
CDE (ours) 0.29 ± 0.35 0.27 ± 0.13
Table 2: Performance comparison of different algorithms trained using our custom curriculum without rehearsal. The score (lower is better $\downarrow$) and completion rate (higher is better $\uparrow$) are reported. The best score and completion rate are highlighted in bold. We see that our proposed method CDE achieves the highest performance and benefits the most from the custom curriculum.
Conclusion
We presented a curriculum for TSP and a continual expansion approach to further improve the performance. Our curriculum design philosophy emphasizes a comprehensive approach to agent training, addressing key skills to enhance generalization to newer environments. Our results show that curriculum design should transcend increasing the size of the environment and instead be tailored to the specific application domain. Fully leveraging such curricula requires a careful balance between retaining previous skills and acquiring new ones, ensuring that agents can adapt to dynamic challenges without falling into suboptimal behaviors or parameter spaces. Our Continual DQN Expansion approach effectively manages the interplay between plasticity and stability. By demonstrating superior improvements and minimal forgetting, our method addresses the limitations of previous approaches. Our results highlight the importance of a holistic curriculum and algorithm design for improved generalization. Thus, there is a need for ongoing research to scale these approaches across different domains for more robust and domain specific continual learning agents.